OpenStudy (anonymous):

Help! Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4). y = four thirdsx + 8 y = four thirdsx + 3 y = -three fourthsx + 8 y = -three fourthsx + 3

OpenStudy (perl):

hint: For the given equation , first solve for y

OpenStudy (anonymous):

how do i solve for y? sorry i dont know how to do this...

OpenStudy (anonymous):

@perl

OpenStudy (perl):

$$ \Large {4x - 3y = -12\\\\ \\ \iff \text{subtract 4x from both sides}\\~\\ -3y = -4x - 12 \\\\ \iff \text{divide both sides by -3}\\~\\ y = \frac{-4}{3}x - \frac{12}{3} } $$

OpenStudy (anonymous):

y=1.3x - 4?

OpenStudy (anonymous):

im not sure what the next step is after what you posted

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

i need to go to sleep, its 2:10 AM and i have to wake up at 5:20 for school :( can some one please help me?

OpenStudy (butterflydreamer):

perl already walked you through it :) what don't you understand?

OpenStudy (anonymous):

what do i have to do after y=-4/3x-12/3

OpenStudy (butterflydreamer):

actually firstly i think... there may be an error in that equation :) let's start over.

OpenStudy (anonymous):

alright :)

OpenStudy (butterflydreamer):